Solve the question one in exercise one
You stand on the edge of a canyon 420 m and clap your hands; the echo comes back to you 2.56 sec later. What is the temperature of the air? (γ_air=1.40 and M=28.8gm)
Solution
The speed of sound in a gas is v=√(β/ρ)=√(γRT/M)
⇒T=(Mv^2)/γR=[(2.88×10^(-2) kg)(328 m/s^2 )]⁄([(1.4)(8.31)]=266K=-7℃)
1
Expert's answer
2020-01-03T09:18:50-0500

First of all let's calculate the velocity of the sound wave"=2*d\/t"

where distance is taken as (d+d) because sound wave travels from speaker to canyon and then it returns back to the speaker and t is the time taken by the sound wave to cover the (d+d) distance

velocity of sound wave "=" "2*420\/2.56=328m\/s"

velocity of the sound wave ="\sqrt{\smash[b]{{\gamma}RT/M}}"

where "\gamma" "=1.4" , "R""=8.314", and "M=28.8*10^{-3}kg"

putting these in the above equation ,

"328= \sqrt{\smash[b]{{1.4*8.314*T/(28.8*10^{-3})}}}"

on calculating we get,

"T=266.2K"

In "^0C" ,

"T" "=266.2-273=-7.2^0C"